# Causal Lifting of Neural Representations: Zero-Shot Generalization for Causal Inferences

### v2

@@ Abstract
In many scientific domains, the cost of data annotation limits the scale and pace of experimentation. Yet, modern machine learning systems offer a promising alternative—provided their predictions yield correct conclusions. We focus on Prediction-Powered Causal Inferences (PPCI), i.e., estimating the treatment effect in a target experiment with unlabeled factual outcomes, retrievable zero-shot from a pre-trained model. We first identify the conditional calibration property to guarantee valid PPCI at population level. Then, we introduce “causal lifting”, a new causal lifting constraint transferring validity across experiments, which we propose to enforce in practice in Deconfounded Empirical Risk Minimization, our new model-agnostic training objective. We validate our method on synthetic and real-world scientific data, offering solutions to instances not solvable by vanilla Empirical Risk Minimization and invariant training. In particular, we solve zero-shot PPCI on the ISTAnt dataset for the first time, fine-tuning a foundational model on our replica dataset of their ecological experiment with a different recording platform and treatment.
@@ 1 Introduction
Artificial Intelligence (AI) systems hold great promise for accelerating scientific discovery by providing flexible models capable of automating complex tasks. We already depend on deep learning predictions across various scientific domains, including biology (Jumper et al., 2021; Tunyasuvunakool et al., 2021), medicine (Elmarakeby et al., 2021; Mullowney et al., 2023), sustainability (Zhong et al., 2020; Rolnick et al., 2022), and social sciences (Jerzak et al., 2022; Daoud et al., 2023). While these models offer transformative potential for scientific research, their black-box nature poses new challenges, especially when used to analyze new experimental data. They can perpetuate hidden biases, which are difficult to detect and quantify, potentially invalidating any analysis resting on their predictions.
Recent works proposed leveraging predictive models to offer valid and efficient statistical inferences on partially labeled data (Angelopoulos et al., 2023a, b), using the labeled data to adapt the model and then estimating the statistical quantities of interest using the model’s predictions on the unlabeled samples. In this paper, we focus on powering causal inferences on fully unlabeled experimental data, where an outcome of interest is captured in high-dimensional measurements such as images or videos. Together with other experimental and treatment information, the outcome predictions can be used to quantify the interventional effect of the treatment. We refer to this problem as zero-shot Prediction-Powered Causal Inference (PPCI). A key challenge is that small modeling biases can invalidate the causal conclusions, even in the simplest possible scenario, where the target experiment is a randomized controlled trial that is partially annotated (Cadei et al., 2024), so one can train on it. Still, being able to solve a PPCI problem zero-shot, i.e., without training on the data from the target experiment, would significantly accelerate scientific discovery, for example, providing the conclusion of a new experiment in real time, without any costly and time-consuming human annotations.
To tackle this problem, we first characterize when a predictor is valid for a given PPCI problem. We show that (statistical) conditional calibration with respect to suitable experimental settings transfers the identification and estimation properties of the AIPW estimator (Robins et al., 1994) to the prediction-powered samples. This result considers an arbitrary predictor, thus opening the opportunity of training on some other similar experiment that shares the same outcome-measurement mechanism (i.e., concerns the same outcomes and collects the same input modality), so that it can be applied zero-shot to the new experimental data. Unfortunately, vanilla training on a similar experiment with empirical risk minimization does not guarantee valid predictors. Particularly, spurious associations between the training and target experiments may not be resolved, in some cases, even with infinite training data. To fix this, we propose Causal Lifting, a representation learning constraint that we show transfers causal validity under standard (yet idealized) assumptions.
@@
Our experimental results are inspired by Cadei et al. (2024), which is to date the only real-world representation learning benchmark for treatment effect estimation. Tailoring to the discrete treatment assignments and experimental settings recorded in this benchmark, we propose a simple yet effective implementation of the causal lifting constraint via loss reweighting, which we call Deconfounded Empirical Risk Minimization (DERM). To showcase zero-shot generalization to the ISTAnt dataset, we create and release a new dataset, from our similar ecological experiment with a different recording platform and treatments. Remarkably, although our replica dataset is of lower quality compared to the ISTAnt experiment, we show for the first time valid PPCI on ISTAnt, which we surprisingly achieve zero-shot. We further validate and confirm the results on a synthetic manipulation of MNIST dataset (LeCun, 1998) that allows controlling for the true causal effect.
@@
    a new formulation and analysis for Prediction-Powered Causal Inference, preserving estimation properties, motivated by real-world scientific applications,
@@
    Causal Lifting, a new constraint to transfer causal validity zero-shot across PPCI problems, alongside a practical implementation (DERM),
@@
    first valid 0-shot prediction-powered Average Treatment Effect (ATE) inference on ISTAnt (Cadei et al., 2024), leveraging a new dataset collected from our similar experiment.
@@ 2 Prediction-Powered Causal Inference
Problem: Prediction-Powered Causal Inference (PPCI) Let $(T,W,Y,X)\sim\mathcal{P}$ be random variables, where $T\in\mathcal{T}$ denotes a treatment assignment, $W\in\mathcal{W}$ the observed covariates, $Y\in\mathcal{Y}$ the outcome of interest, and $X\in\mathcal{X}$ an high-dimensional measurement used to predict $Y$. Let $g:\mathcal{X}\to\mathcal{Y}$ be a fixed predictor of $Y$ from $X$. Given an i.i.d. sample $\{(T_{i},W_{i},\_,X_{i})\}_{i=1}^{n}\overset{i.i.d.}{\sim}\mathcal{P}$, where the outcomes are not observed directly, identify and estimate the average potential outcome under a treatment intervention [^1] We focus on population-level causal estimands, but the problem generalizes to group-level quantities too by additionally conditioning on a covariates subset., i.e., $\tau_{Y}(t):=\mathbb{E}[Y\mid\operatorname{do}(T=t)]$, or related estimands, from the prediction-powered sample $\{(T_{i},W_{i},g(X_{i}))\}_{i\in\mathcal{I}}$.
@@
Several scientific problems fit under this framework, where a human domain expert, trained on similar data, acts as a predictor $g$, elaborating and annotating the high-dimensional measurements, aiming to collect evidence to answer the overarching scientific question. Replacing this step with a deep learning model requires strong guarantees, which generally differ from standalone prediction accuracy metrics. Systematic errors in a specific subgroup can invalidate the full analysis, as shown in Cadei et al. (2024). However, the PPCI potential comes from the fact that good outcome models can be learned also from similar experiments and transferred zero-shot, e.g., changing set-up and treatment of interest, as far as the (true) annotation mechanism stays invariant. Such a generalization goal is the main motivation behind the investigation of this problem, going beyond (statistical) Prediction-Powered Inference (Angelopoulos et al., 2023a, b) that is designed instead for efficiency in-distribution. As a practical, real-world motivating example, we propose to solve for the first time ATE estimation on the ISTAnt dataset zero-shot, i.e., without behaviour annotations. Instead, we leverage our own experiment variant with different treatment and recording platform, see Figure 1.
@@
Figure 1: Illustration of our PPCI pipeline for (zero-shot) ATE estimation on ISTAnt dataset. We designed, recorded and annotated our own ISTAnt version with different treatments and recording platform and leveraged it to fine-tune a pre-trained model for imputation.
@@
Before discussing the generalization challenges and our proposed mitigation, we formalize the properties a predictor needs to support valid PPCI and what the guarantees are. For this, note that, in the following discussion, we always start from assuming identifiability of the average potential outcome(s). Indeed, if the problem has no solution even with access to the true outcomes, the corresponding PPCI problem loses meaning.
@@
An outcome predictor $g$ is valid for a PPCI problem if and only if the average potential outcome(s) of interest $\tau_{Y}(t)$ identify in a statistical estimand, i.e., a treatment function removing the do-operator, and $\tau_{Y}(t)$=$\tau_{g(X)}(t)$.
For example, given a PPCI problem $\mathscr{P}$ with binary treatment, let the average potential outcome upon intervening on the treatment value $t$ being identified via the adjustment formula by standard causal inference assumptions [^2] Consistency, exchangeability and overlap., i.e.,
@@
where $W$ is a valid adjustment set. A predictor $g$ is valid for $\mathscr{P}$ if $\tau_{Y}(t)$=$\tau_{g(X)}(t)$, i.e.,
@@
Then, we want to characterize testable statistical property a model has to satisfy in order to be valid.
@@
The lemma follows directly from backdoor adjustment (see Proof in Appendix A), and offer a sufficient condition for validity. It is worth observing that being calibrated with respect to all the observed covariates and not only the valid adjustment set is just a stronger sufficient condition, still holding the thesis. For example, in experiments where there are few discrete experimental conditions, it is safe to enforce/test it for all the covariates, without requiring a prior knowledge of which is the valid adjustment set [^3] Note that the requirements here are different than covariate adjustment, and calibrating only to a minimal set is not necessary if the model can be practically calibrated on more variables without loosing predictive performances.. {theorembox}
@@
Given a PPCI problem $\mathscr{P}$ with standard causal identification assumptions. If an outcome model $g:\mathcal{X}\rightarrow\mathcal{Y}$ is conditionally calibrated with respect to the treatment and a valid adjustment set, then the AIPW estimator  (Robins et al., 1994) over the prediction-powered sample preserves the asymptotically valid confidence interval for the true ATE, i.e.,
@@
As before, the result is intuitive and now extend such sufficient condition to preserve the property of an estimator, i.e., asymptotic normality. The technical proof follows from classical AIPW convergence results, and we report it in Appendix A. More generally about estimation, it is important to observe that, while several estimators may exist for the same causal estimand, it is crucial to rely on a valid predictor for the statistical estimand we identify the causal one with. For example, in a randomized controlled trial, we may identify the ATE with the associational difference, and therefore require calibration only with respect to the treatment; if in practice we want to use an estimator relying on covariates conditioning to improve efficiency, e.g., AIPW, we must require the outcome model to condition on all such covariates too. We still have not discussed how to enforce or even test such a property, not having access to outcome annotations on the target experiment. For this, we need to discuss first the crucial generalization challenges in learning a valid outcome model.
@@ 2.1 Out-of-distribution generalization challenges
The main challenge in PPCI is training a valid outcome model to replace expensive and slow annotation procedures. It has to be expressive enough to process the complex input measurements, while being calibrated with respect to the experimental settings. Despite modern pre-trained models are getting more and more capable and generalist (Bubeck et al., 2024), they still cannot guarantee calibration a-priori (Chen et al., 2022). As a human annotator learns an annotation procedure from similar experiments with the same measurement type (e.g., videos) and outcome of interest (e.g., ants behaviors), we may aim to fine-tune pre-trained models on similar annotated experiments, leading towards scientifically valid expert models. For this purpose, let’s first provide a notion of similarity among experiments.
@@
This definition naturally generalizes to whatever measurement-outcome couple $(X,Y)\sim\mathcal{P}$ such that the annotation mechanism $\mathbb{P}(Y|X=x)$ is invariant. Note that our setting is different from the invariant causal predictions (Peters et al., 2016), as in our case $\mathbb{P}(Y|X=x)$ is not a causal mechanism that we want to identify using an invariance assumption over multiple environments, but rather $X$ is a measurement of $Y$ (Silva et al., 2006) (and thus a child of Y), potentially observed over a single environment $\mathscr{P}$. We can rely on this anticausal invariance whenever the same annotation protocol could be used for the two experiments (for example, humans trained to annotate $\mathscr{P}$ are capable of annotating $\mathscr{P}^{\prime}$ without new training or instructions). Still, vanilla fine-tuning by standard empirical risk minimization over similar experiments doesn’t guarantee learning a valid outcome model due to peculiar generalization challenges raised by the spurious shortcuts appearing in anticausal prediction problems (Schölkopf et al., 2012; Peters et al., 2016), for example, from the visual appearance of certain experimental settings. We distinguish here between statistical challenges with finite training samples and potential issues even in the infinite sample regime.
@@
For a certain training subsample the model may over-rely on its spurious conditional measurement-outcome correlation, e.g., some irrelevant background information. If on target the measurements in such subgroup are over-representing a certain treatment value or other outcome’s ancestors, it may invalidate any causal downstream prediction-powered estimation.
Example 1: Background confounding [Uncaptioned image] Consider PPCI within the ISTAnt dataset, annotating a first batch of videos for training. If a certain behaviour is underrepresented in a specific plate, e.g., because the plate is in the control group, the model can leverage such spurious shortcut in subsequent batches, relying on background information like the pen markings on the floor to identify the plate. This invalidates any downstream prediction-powered analysis, as future plates may be treated, but the model may assume they are not, matching the target statistics with the control group.
@@
Additionally, treatment information captured in the measurement that is not mediated by the outcome may be misleading for generalization, even in the infinite training sample setting. In other words, a double blind experiment needs to be blind also for the model, which should not be able to guess the treatment if not through the effect. Otherwise, prior correlations between other treatments and the effect can propagate to the new experiment. For example, say a model was trained on an experiment using treatment $T^{\prime}=t^{\prime}$. In the target experiment, we consider a different treatment $T=t$ that “looks like $t^{\prime}$”, i.e., given an image that may come from either experiment, $t$ and $t^{\prime}$ are equally likely:
$\mathbb{P}(T=t|X=x)=\mathbb{P}(T^{\prime}=t^{\prime}|X^{\prime}=x)\quad\forall x\in supp(X)\cap supp(X^{\prime}).$ (6)
The model, that is only trained on $t^{\prime}$ and has never seen $t$ during training, may introduce correlations with the training treatment effect. For example, if an outcome value never happens in the training treatment ($supp(Y^{\prime}|T^{\prime}=t^{\prime})\subset supp(Y^{\prime})$), the model will wrongly transfer this to the target population, invalidating any causal downstream prediction-powered estimation. The same argument also holds replacing the treatment with any other outcome’s ancestor.
Example 2: Different treatments with similar appearance [Uncaptioned image] Consider two experiments where the treatments are visually similar, discernible from the control, and differing in the effect, e.g., different pathogen suspension exposure with liquid residual in the measurements. While learning, an outcome model may over-rely on the training treatment effect, e.g., excluding the possibility of a specific behaviour, and wrongly transfer during PPCI on the target experiment. Note that this is not the case for our experiment and ISTAnt where the treatment is not visible, as for well-designed double blind experiments.

### v3

# Prediction-Powered Causal Inferences

@@ Abstract
In many scientific experiments, the data annotating cost constraints the pace for testing novel hypotheses. Yet, modern machine learning pipelines offer a promising solution—provided their predictions yield correct conclusions. We focus on Prediction-Powered Causal Inferences (PPCI), i.e., estimating the treatment effect in an unlabeled target experiment, relying on training data with the same outcome annotated but potentially different treatment or effect modifiers. We first show that conditional calibration guarantees valid PPCI at population level. Then, we introduce a sufficient representation constraint transferring validity across experiments, which we propose to enforce in practice in Deconfounded Empirical Risk Minimization, our new model-agnostic training objective. We validate our method on synthetic and real-world scientific data, solving impossible problem instances for Empirical Risk Minimization even with standard invariance constraints. In particular, for the first time, we achieve valid causal inference on a scientific experiment with complex recording and no human annotations, fine-tuning a foundational model on our similar annotated experiment.
@@ 1 Introduction
Artificial Intelligence systems hold great promise for accelerating scientific discovery by providing flexible models capable of automating complex tasks. We already depend on deep learning predictions across various scientific domains, including biology (Jumper et al., 2021; Tunyasuvunakool et al., 2021), medicine (Elmarakeby et al., 2021; Mullowney et al., 2023), sustainability (Zhong et al., 2020; Rolnick et al., 2022), and social sciences (Jerzak et al., 2022; Daoud et al., 2023). While these models offer transformative potential for scientific research, their black-box nature poses new challenges, especially when used to analyze new experimental data. They can perpetuate hidden biases, which are difficult to detect and quantify, potentially invalidating any analysis resting on their predictions (Cadei et al., 2024).
Recent works proposed leveraging predictive models for efficient and still valid statistical inferences on partially labeled data (Angelopoulos et al., 2023a, b), training the predictor on the annotated sample and then estimating the statistical quantities of interest, including the model’s predictions on the unlabeled sample. In this paper, we focus on causal inferences from (outcome) unlabeled experiments. We similarly aim to retrieve the missing labels from their complex measurements, i.e., entangled and high-dimensional, such as images or videos, through a predictive model. However, we consider training such a model from similar yet different annotated experiments, e.g., different treatments or effect modifiers. We refer to this problem as Prediction-Powered Causal Inference (PPCI), differing from classic Prediction-Powered Inference for its motivation, i.e., causal, and setting, i.e., generalization.
To tackle this problem, we first characterize when a predictor is valid for a given PPCI problem. We show that (statistical) conditional calibration with respect to the observed outcome parents preserves identifiability and still guarantees consistent estimation, e.g., via Augmented Inverse Propensity Weighted (AIPW) estimator (Robins et al., 1994), for the treatment effect in the unlabeled experiment. However, the validity of a predictor may not transfer even between similar experiments and assuming an infinite training sample by Empirical Risk Minimization, even with standard invariance constraints (Arjovsky et al., 2019; Krueger et al., 2021). We then propose a sufficient condition to enforce in the representation backbone of the predictor, causal lifting the model to transfer causal validity on the target experiment (together with other standard, yet idealized, assumptions).
@@
We further propose a novel, simple yet effective implementation of such a constraint via loss reweighting, which we name Deconfounded Empirical Risk Minimization (DERM). To test it, we considered ISTAnt experiment (Cadei et al., 2024) (unique real-world benchmark for treatment effect estimation with complex measurements), ignoring the outcome annotations, and trained the predictive model over a new annotated experiment of ours with the same annotation mechanism, but different recording platform (lower quality) and treatments. We further validate and confirm the results on a synthetic manipulation of MNIST dataset (LeCun, 1998) by controlling the data-generating process and causal effect.
@@
    formulation and foundational theory for Prediction-Powered Causal Inferences, i.e., prediction-powered causal estimands identification and estimators properties,
@@
    a new constraint to transfer causal validity across similar experiments, causal lifting neural representations with out-of-distribution fine-tuning, alongside a practical implementation (DERM),
@@
    first valid causal inference on a scientific experiment (ISTAnt) with complex measurements and no human annotations, by fine-tuning a model on our similar annotated experiment.
@@ 2 Prediction-Powered Causal Inference
Problem: Prediction-Powered Causal Inference (PPCI) Let $(T,W,Y,X)\sim\mathcal{P}$ be random variables, where $T\in\mathcal{T}$ denotes a treatment assignment, $W\in\mathcal{W}$ the observed covariates, $Y\in\mathcal{Y}$ the outcome of interest, and $X\in\mathcal{X}$ a complex measurement capturing $Y$ information. Let $g:\mathcal{X}\to\mathcal{Y}$ be a fixed predictor of $Y$ from $X$. Given an i.i.d. sample $\{(T_{i},W_{i},\_,X_{i})\}_{i=1}^{n}\overset{i.i.d.}{\sim}\mathcal{P}$, where the outcomes are not observed directly, identify and estimate the average potential outcome under a treatment intervention [^1] We focus on population-level causal estimands, but the problem generalizes to group-level quantities too by additionally conditioning on a covariates subset., i.e., $\tau_{Y}(t):=\mathbb{E}[Y\mid\operatorname{do}(T=t)]$, or related estimands, from the prediction-powered sample $\{(T_{i},W_{i},g(X_{i}))\}_{i=1}^{n}$.
@@
Several scientific problems fit under this framework, where a human domain expert, trained on similar data, acts as a predictor $g$, elaborating and annotating the high-dimensional measurements, aiming to collect evidence to answer the overarching scientific question. Replacing this step with a deep learning model requires guarantees, which generally differ from standalone prediction accuracy metrics. Systematic errors in a specific subgroup can invalidate any causal reasoning (Cadei et al., 2024; Cadei and Internò, 2025). However, the PPCI potential comes from the fact that good outcome models can be learned also from similar experiments and transferred zero-shot, e.g., changing setup and treatment of interest, as long as the (true) annotation mechanism stays invariant. Such a generalization goal is the main motivation behind the investigation of this problem, going beyond (statistical) Prediction-Powered Inference (Angelopoulos et al., 2023a, b) that is designed instead for efficiency in-distribution. As a practical, real-world motivating problem, we aim to achieve valid Average Treatment Effect estimation on ISTAnt without expensive human annotations, but by annotating instead with foundational models fine-tuned over our annotated similar experiment, with different treatments and recording platform. See the problem illustration in Figure 1.
@@
Figure 1: Pipeline for Average Treatment Effect inference on ISTAnt experiment without human annotations. We designed, recorded, and annotated our similar experiment with different treatments and recording platform and leveraged it to fine-tune a pre-trained model for imputation.
@@
Before discussing the generalization challenges and our proposed mitigation, we formalize the properties a predictor requires to guarantee valid PPCI. Note that, in the following discussions, we always start by assuming identifiability of the average potential outcome(s), e.g., Randomized Controlled Trial. Indeed, if the problem has no solution even with access to the true outcomes, the corresponding PPCI problem loses meaning  [^2] Unless an unobserved and valid adjustment set can be identified from the complex measurements $X$, but such a discussion is outside the scope of this paper..
@@
An outcome predictor $g$ is valid for a PPCI problem if and only if the average potential outcome(s) of interest $\tau_{Y}(t)$ identify in a statistical estimand, i.e., a treatment function removing the do-operator, and $\tau_{Y}(t)$$\tau_{g(X)}(t)$.
For example, given a PPCI problem $\mathscr{P}$ with binary treatment, let the average potential outcome upon intervening on the treatment value $t$ be identified via the adjustment formula by standard causal inference assumptions [^3] Consistency, exchangeability, and overlap., i.e.,
@@
where $W$ is a valid adjustment set. A predictor $g$ is valid for $\mathscr{P}$ if $\tau_{Y}(t)$$\tau_{g(X)}(t)$, i.e.,
@@
Then, we want to characterize the testable statistical property a model has to satisfy to be valid.
@@
Given a PPCI problem $\mathscr{P}$ with identifiable ATE (given the ground-truth annotations). If an outcome model $g:\mathcal{X}\rightarrow\mathcal{Y}$ is conditionally calibrated with respect to the treatment and the valid adjustment set considered for identification, then it is (causally) valid for $\mathscr{P}$.
The lemma follows directly from backdoor adjustment, see proof in Appendix A, and offers a sufficient condition for validity. It is worth observing that being calibrated with respect to all the observed covariates and not only the valid adjustment set is just a stronger sufficient condition, still holding the thesis. For example, in experiments where there are few discrete experimental conditions, it is safe to enforce/test it for all the covariates, without requiring a prior knowledge of which is the valid adjustment set.{theorembox}
@@
Given a PPCI problem $\mathscr{P}$, with identifiable ATE (given the ground-truth annotations). If an outcome model $g:\mathcal{X}\rightarrow\mathcal{Y}$ is conditionally calibrated with respect to the treatment and a valid adjustment set, then the corresponding AIPW estimator over the prediction-powered sample preserves doubly-robustness and valid asymptotic confidence intervals, i.e.,
@@
As before, the result is intuitive and now extends such a sufficient condition to preserve the property of an estimator. The proof, despite being technical, follows directly from classical AIPW results, and we present it in Appendix A. In practice, it is important to note that, while several estimators may exist for the same causal estimand, we need to rely on a valid predictor for the statistical estimand we identify the causal one with. For example, in a randomized controlled trial, we may identify the ATE with the associational difference, and therefore require calibration only with respect to the treatment; if, in practice, we want to use an estimator relying on covariates conditioning to improve efficiency, e.g., AIPW, we must require the outcome model to condition on all such covariates too. We still have not discussed how to enforce or even test such a property, without access to outcome annotations on the target experiment, and we start illustrating the fundamental generalization challenges in learning a valid outcome model.
@@ 2.1 Out-of-distribution generalization challenges
The main challenge in PPCI is training a valid outcome model to replace expensive and slow annotation procedures. It has to be expressive enough to process the complex input measurements, while being calibrated with respect to the experimental settings. Although modern pre-trained models are getting more and more capable and generalist (Bubeck et al., 2024), they still cannot guarantee calibration a priori (Chen et al., 2022). As a human annotator learns an annotation procedure from similar experiments with the same measurement type (e.g., videos) and outcome of interest (e.g., ants’ behaviours), we may aim to fine-tune pre-trained models on similar annotated experiments, leading towards scientifically valid expert models. For this purpose, let’s first provide a notion of similarity among experiments.
@@
This definition naturally generalizes to whatever measurement-outcome couple $(X,Y)\sim\mathcal{P}$ such that the annotation mechanism $\mathbb{P}(Y|X=x)$ is invariant. Note that our setting is different from the invariant causal predictions (Peters et al., 2016), as in our case $\mathbb{P}(Y|X=x)$ is not a causal mechanism that we want to identify using an invariance assumption over multiple environments, but rather $X$ is a measurement of $Y$ (Silva et al., 2006; Yao et al., 2025) (and thus a child of Y), potentially observed over a single environment $\mathscr{P}$. We can rely on this anticausal invariance whenever the same annotation protocol could be used for the two experiments (for example, humans trained to annotate $\mathscr{P}$ are capable of annotating $\mathscr{P}^{\prime}$ without new training or instructions). Still, vanilla fine-tuning by standard Empirical Risk Minimization over similar experiments doesn’t guarantee learning a valid outcome model due to peculiar generalization challenges raised by the spurious shortcuts appearing in anticausal prediction problems (Schölkopf et al., 2012; Peters et al., 2016), for example, from the visual appearance of certain effect modifiers. We distinguish here between statistical challenges with a finite training sample and potential issues even in the infinite sample regime.
@@
If the training sample is not representative of the full target population, the model may over-rely on its spurious conditional measurement-outcome correlation, e.g., some irrelevant background information. Concretely, the measurements in such a subgroup are over-representing a certain treatment value or effect modifiers, e.g., up to violating positivity, which may bias any training and invalidate any causal downstream prediction-powered estimation.
Example 1: Background confounding [Uncaptioned image] Consider PPCI for ISTAnt, annotating only the first batch of nine videos for training. If a certain behaviour is underrepresented in a specific plate, e.g., because the plate is in the control group, the model can leverage such spurious shortcut in subsequent batches, relying on background information like the pen markings on the floor to identify the plate. This invalidates any downstream prediction-powered analysis, as future plates may be treated, but the model may assume they are not, matching the target statistics with the control group.
@@
Additionally, treatment information captured in the measurement that is not mediated by the outcome may be misleading for generalization, even in the infinite training sample setting. In other words, a double blind experiment needs to be blind also for the model, which should not be able to guess the treatment and other effect modifiers if not through the effect. Otherwise, the causal relationships in the training data can be wrongly transferred to the target experiment, potentially on different variables and distributions. For example, if an outcome value never happens in the training treatment, i.e., $supp(Y^{\prime}|T^{\prime}=t^{\prime})\subset supp(Y^{\prime})$, the model will wrongly transfer this information to a target population with a different treatment but with the same appearance, invalidating any causal downstream prediction-powered estimation there.
Example 2: Different treatments with similar appearance [Uncaptioned image] Consider two experiments where the treatments are visually identical, discernible from the control, and differing in the effect, e.g., different pathogen suspension exposure with liquid residual in the measurements. While learning, an outcome model may over-rely on the training treatment effect, e.g., excluding the possibility of a specific behaviour, and wrongly transfer during PPCI on the target experiment. Note that this is not the case for our experiment and ISTAnt where the treatment is not visible, as for well-designed double blind experiments.